OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 7 is 6 less than –8. What is the number? A. –98 B. –56 C. –14 D. –7

OpenStudy (anwaarullah):

C -14

OpenStudy (anwaarullah):

x/7= 6-8 x=-2*7